Circular No. 7425

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cj IN NGC 6753 Robert Evans, Hazelbrook, N.S.W., reports the visual discovery of an

apparent supernova (mag ∼ 14.8) on May 14.7 UT with a 0.31-m reflector. A CCD image taken in bright moonlight by T. Dobosz, Bankstown, N.S.W., with a 0.28-m Schmidt-Cassegrain telescope on May 15.6 shows the new star superimposed upon a spiral arm of the galaxy, and a measurement of this image by G. L. White, Nepean Centre for Astronomy, Kingswood, N.S.W., yields the following position for SN 2000cj: α = 19^{h}11^{m}27^{s}.49, δ = −57^{o}03′14′′.1 (equinox 2000.0), which is ∼ 35′′ east and 19′′ south of the nucleus of NGC 6753. Position end figures for SN 2000cj (mag 14.7) measured by R. H. McNaught at Siding Spring (from a CCD image obtained by R. Langersek, Macgregor, Qld., with a 0.25-m Schmidt-Cassegrain telescope on May 15.62) are 27^{s}.68, 19′′.4 (reduction using ten USNO-A2.0 stars, rms 0′′.4), with the offset being 33′′ east and 22′′ south of the galaxy center. McNaught confirms that nothing appears at this position on the ESO B, ESO R, and SERC J surveys. N. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), reports that preliminary photometry from observations obtained by L. Strolger on May 17.1 with the CTIO 0.9-m telescope yields V = 14.56, B − V = 0.00. Additional visual magnitude estimates: May 14.79, 14.8: (P. Marples, Loganholme, Qld.); 15.7, 14.7 (Evans).

LMC X-3 J. Homan, P. G. Jonker, and M. van der Klis, Astronomical Insti-

tute ‘Anton Pannekoek’, University of Amsterdam; and E. Kuulkers, Space Research Organization Netherlands, Utrecht, and Astronomical Institute, Utrecht University, report: “The blackhole candidate LMC X-3 was observed with the RXTE/PCA during May 13.894–13.995 UT. The 2–30-keV power spectrum of that observation reveals strong variability, with a fractional rms amplitude of ∼ 40 percent (0.01–100 Hz). The shape of the power spectrum can be characterized by a broken power law, with a break at 1.1 ± 0.4 Hz. In addition a quasiperiodic oscillation was found at 0.42 ± 0.02 Hz, with a strength of 12 ± 2 percent (3.7σ, single trial), and a width of 0.10 ± 0.04 Hz. Spectral fits yield values similar to those reported by Boyd and Smale (IAUC 7424). This is the first time that LMC X-3 has been found to show rapid time variability similar to that of blackhole candidates in their low/hard state. Together with the spectral measurements, this strongly suggests that LMC X-3 has entered such a low/hard state.”

2000 May 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green